OpenStudy (anonymous):

Simplify the result, and leave your answer in factored form. (x/x^2 - 3x + 2) - (x/x^2 - x - 2) The answer I got was x/(x - 1)(x - 2)(x + 1), so could someone tell me if it's right, thanks!

OpenStudy (anonymous):

some mistake here denominator is right, but i think the numerator should be \(2x\)

OpenStudy (anonymous):

\[\frac{x}{(x-1)(x-2)}-\frac{x}{(x+1)(x-2)}\] \[\frac{x(x+1)-x(x-1)}{(x-1)(x+1)(x-2)}\] \[\frac{x^2+x-x^2+x}{(x-1)(x+1)(x-2)}\]

OpenStudy (anonymous):

Yeah I got that far, but how does the top turn into 2x? I though the x^2 and the x would cancel out or something, idk. o;

OpenStudy (anonymous):

\[x^2+x-x^2+x=\cancel{x^2}+x-\cancel{x^2}+x=x+x=2x\]

OpenStudy (anonymous):

you did all the hard part. combining like term should be easy part

OpenStudy (anonymous):

Ohhh. I got a sign wrong, thanks!

OpenStudy (anonymous):

it is that damned distributive law, gets you every time

OpenStudy (anonymous):

It sure does Lool.